OpenStudy (maleena_1):

A conveyor belt carries supplies from the first floor to the second floor, which is 12 feet higher. The belt makes a 60 degree angle with the ground. How far do the supplies travel from one end of the conveyor belt to the other? Round your answer to the nearest foot.

563blackghost (563blackghost):

Which Trigonometry Identity would we use based on this visual?

563blackghost (563blackghost):

@maleena_1

OpenStudy (maleena_1):

cos

OpenStudy (maleena_1):

@563blackghost

563blackghost (563blackghost):

Actually we would use sin becuase we know of the opposite and want to know of the hypotenuse side.... Now we insert this into the ratio of sin...

OpenStudy (maleena_1):

I don't get the ratio part

563blackghost (563blackghost):

\(\huge\color{purple}{\sin(60)=\frac{12}{x}}\) We would divide by 12 since 12 is located as the numerator and we need the variable x as its self... \(\huge\color{red}{x=12\div\sin(60)}\) Note: When inputting this into your calculator make sure that your calculator is set to degree instead of radian mode....

563blackghost (563blackghost):

\(\Large{\cos(x)=\frac{adjacent}{hypotenuse}}\) \(\Large{\sin(x)=\frac{opposite}{hypotenuse}}\) \(\Large{\tan(x)=\frac{opposite}{adjacent}}\) These are the three identities and their ratios...

OpenStudy (maleena_1):

I got a decimal? 0.866

563blackghost (563blackghost):

That is incorrect.... did you place the problem in the calculator as shown... When inputting it would look like.... \(\Large{12/\sin(60)}\)

OpenStudy (maleena_1):

yes that is what I put in

563blackghost (563blackghost):

And is your calculator set in degree mode?....to check look on you calculator for mode; then you would click it and find "radian" and "degree" in which you would click on "degree" and then leave the area...

OpenStudy (maleena_1):

yes it is degree

563blackghost (563blackghost):

Well thats weird... O.o

OpenStudy (maleena_1):

how am I supposed to get the answer

563blackghost (563blackghost):

Well you should get the total of 13.8564... which is the length of the conveyor belt distance... now do you need this in decimal or radical form?

OpenStudy (maleena_1):

I got it thanks

563blackghost (563blackghost):

\(\Huge\cal\rlap{\color{red}{No~Problem!:)}}{\color{blue}{\;No~Problem!:)}}\) I hope your cal will work for future problems :)